If y varies directly as x and y=8 when x=3 find y when x=5
Answer by stanbon(75887)

y varies directly as x means the following:
y= kx where "k" is a number we have to find.
Since y=8 when x=3 we get the following equation
which allows us to solve for "k": 8=k(3)
Therefore k=8/3
Now we know the general formula is y=(8/3)x
So, if x=5 y=(8/3)(5) = 40/3
Cheers,
Stan H.